OpenStudy (anonymous):

I always had trouble doing these..Can someone clearly explain on how to do these type of problems starting with this question: f(x)= 4x/(1-x) and g(x)= 2/x, then f(g(x))=?

OpenStudy (anonymous):

ok. I'll try

OpenStudy (anonymous):

work from the middle of the f(g(x)) g(x)=2/x so everytime you see a g(x), plug in 2/x, because they're the same thing.

OpenStudy (anonymous):

like this f((2/x)) ??

OpenStudy (zarkon):

\[f(g(x))=f\left(\frac{2}{x}\right)\] yes

OpenStudy (anonymous):

Now, help me out Zarkon if I'm wrong here, but I believe u plug in the f(x)equation and put 2 over it

OpenStudy (zarkon):

every where you see an x in the function f(x) replace it with 2/x

OpenStudy (anonymous):

I think that Zarkon might be right. I was just guessing for mine.

OpenStudy (zarkon):

I am ;)

OpenStudy (anonymous):

Okay so I did that Zarkon, we know have 4(2x)/1-2x and also the 2/x

OpenStudy (zarkon):

try again

OpenStudy (zarkon):

you replaced x with 2x not 2/x

OpenStudy (anonymous):

\[4(2/x)\div \left( 1-(2/x) \right)\]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

$$f(x)=\frac{4x}{1-x}$$ $$f\left(\frac{2}{x}\right)=\frac{4\left(\frac{2}{x}\right)}{1-\left(\frac{2}{x}\right)}$$

OpenStudy (anonymous):

Opps sorry, I corrected it now...so how do we solve that

OpenStudy (zarkon):

that is it

OpenStudy (zarkon):

you can simplify if you like

OpenStudy (anonymous):

cancel out the (2/x)

OpenStudy (zarkon):

multiply top and bottom by x

OpenStudy (anonymous):

dat works too

OpenStudy (anonymous):

then what would u have Sarah blu?

OpenStudy (anonymous):

hmm I am getting the answer 8/-1 .. i am not sure i am understanding when you mean multiply top and bottom by x

OpenStudy (zarkon):

$$\frac{4\left(\frac{2}{x}\right)}{1-\left(\frac{2}{x}\right)}\frac{x}{x}=\frac{8}{x-2}$$

OpenStudy (zarkon):

understand?

OpenStudy (anonymous):

why is the x still in the denominator?

OpenStudy (zarkon):

\[\left(1-\frac{2}{x}\right)x=1\cdot x-\frac{2}{x}\cdot x=x-2\]

OpenStudy (zarkon):

distribute

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

Ohhhhhh that makes soo much sense now..Thank you so much :) !! you too gandalfwiz

OpenStudy (anonymous):

aww thanks!